OpenStudy (anonymous):

∑(1to∞)=(k^2+3k+2)/(x^2+x^k+2) use the ratio test to prove that the series converges for x>1

OpenStudy (anonymous):

can we first factor the top and bottom to get a simplefied form?

OpenStudy (anonymous):

the top can be factored as (k+1)(k+2) but the bottom i cannot factor out

OpenStudy (anonymous):

Is this not a power series? The ratio test says: \[\lim_{k \rightarrow \infty} \left| \frac{((k+1)+1)((k+1)+2)}{x^2+x^{k+1}+2}*\frac{x^2+x^k+2}{(k+1)(k+2)}\right|<1\] Now take the limit and it should fall out :P

OpenStudy (anonymous):

yeah it is not a power series . my only concern was the term with the x^k it wouldn't cancel...

OpenStudy (anonymous):

It definitely is a power series :P You use the ratio test to find the radius of convergence for these.

OpenStudy (anonymous):

Let me get out some paper and see if i can work this limit out :P

OpenStudy (anonymous):

oh. it was given in our instruction that this was not given in a power series. which is why i am confused right now. so if i use the l'hospital's rule to calculate the limit, will the derivative of x^k be k(x)^k-1? or will it be like x^k ln(x)?

OpenStudy (anonymous):

Hmmm...h/o just oneee second.

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

From what I can tell the derivative of x^k would be: ln(k)x^k

OpenStudy (anonymous):

Since you're differentiating with respect to k. However, from there, the limit is exceedingly difficult x.x

OpenStudy (anonymous):

ikr? but i think what i was doing wrong the last two nights working on this problem was i am differentiating the x^k wrong..

OpenStudy (anonymous):

i somehow see some light because of this clarification..:P

OpenStudy (anonymous):

Well if you have: \[\huge\frac{d}{dx}(a^{u(x)})=\frac{du}{dx}\ln(a)a^{u(x)}\] Since you're differentiating with respect to k i think that works out that way. But you end up having to do a product rule and it gets grossss.

OpenStudy (anonymous):

thanks for your help!

OpenStudy (anonymous):

oh thank you very much for that clarification, i somehow calculated the limit to be 1/x which if x >1 the series converges...